# ARS: Automatic Routing Solver with Large Language Models

### v1

@@ 4.2 Analysis
The distribution of problems by the number of constraints is shown in Figure 2. Among the 48 problems, one is a simple VRP without any constraints, while the remaining 47 include one to five constraints. In RoutBench, the distribution reflects the proportions of the full set of 5624 problems. Specifically, problems with three constraints are the most common in the dataset, as three-constraint problems dominate the total number of problems with three or fewer constraints. For more complex problems, those with four and five constraints appear in similar proportions, ensuring a balanced representation of high-complexity scenarios.
@@
Figure 2: Analysis of problem distribution across common problems, RoutBench, and all problems. Among the common problems, one is a simple VRP without any constraints.
@@
5 Experiments
The experiments are conducted on common benchmark problems and RoutBench, which include 48 and 1,000 VRP variants, respectively. These variants cover a wide range of real-world problem types, such as heterogeneous, dynamic, multi-commodity, and others.
To evaluate the performance of these LLM-based methods in handling different VRP variants, we assess their ability to generate correct programs within our solver. To analyze the impact of different solvers on the generated programs, we compare our solver with other solvers (e.g., Gurobi and OR-Tools) to investigate their influence. Additionally, we included experiments with other open-source and proprietary LLMs (e.g., DeepSeek v3 and LLaMA 3.1 70B) to explore their impact on program development.
Figure 3: The frequency of constraint usage is analyzed for both RoutBench-S and RoutBench-H.
Table 3: Comparison of program development performance across different solvers for common problems using standard prompting. The best results among these solvers are highlighted in grey.
Solvers Common Problems
SR $\uparrow$ RER $\downarrow$
Standard Prompting (OR-tools) 12.50% 58.33%
Standard Prompting (Gurobi) 14.58% 41.67%
Standard Prompting (Ours) 41.67% 6.25%
@@ 5.1 Comparison with LLM-Based Methods
The first experiment aims to evaluate the ability of ARS to produce successful programs. We compare our proposed ARS with seven LLM-based methods, including Standard Prompting, Chain of Thought (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), Progressive-Hint Prompting (PHP) (Zheng et al., 2023), Chain-of-Experts (CoE) (Xiao et al., 2023), Self-debug (Chen et al., 2023), and Self-verification (Huang et al., 2024). To focus on program development, all methods are implemented using our backbone algorithm framework to handle various VRP variants.
This experiment is conducted on the common problems and RoutBench. We use two metrics to evaluate the results. The Success Rate (SR) measures the proportion of the program where generated solutions pass the validation process. The Runtime Error Rate (RER) quantifies the percentage of programs that fail due to issues such as runtime errors, incorrect API usage, or syntax mistakes.
As shown in Table 2, ARS significantly outperforms other LLM-based methods in generating correct programs for both common problems and RoutBench. It is evident that these compared methods exhibit unsatisfactory performance across all problems, particularly on the RoutBench-H problems, where their SR is merely around 10%. None of the compared algorithms achieved an SR of 50% or higher. In contrast, ARS achieved an SR of 91.67% on the common problems, successfully addressing 44 out of 48 VRPs. Moreover, on RoutBench, ARS outperformed all seven other LLM-based methods by at least 30% in terms of SR. These results highlight ARS’s competitiveness in addressing complex VRP variants.
@@ 5.2 Comparison with Different Solvers
In the previous experiment, to focus on generating correct programs, all methods use the same solving framework to handle VRP variants. To explore the impact of different solving tools on program development, this experiment uses Standard Prompting with three solving tools (OR-tools, Gurobi, and our backbone) for common problems.
As shown in Table 3, both SR and RER achieve significantly better results when using our framework compared to OR-tools and Gurobi. Under standard prompting, our framework achieves an SR of 41.67%, whereas OR-tools and Gurobi reach only about 10% SR. This is because, in our framework, the LLM only needs to complete the constraint codes in Python, whereas the use of solvers is restricted by their specific implementations and requires highly standardized modeling. For the same VRP variants, our solving framework simplifies program development considerably compared to these solvers. Examples of program codes are provided in the appendix F.
To evaluate the performance of our solver on different VRP variants, we also conducted a comparative study with two other solvers across three common VRP variants (e.g., CVRP, CVRPTW, and CVRPL). The experiments are conducted on instances with 25 and 50 nodes, with time limits of 50 seconds and 5 minutes, respectively.
As can be seen from Table 4, our solver achieved the optimal results in many instances. Notably, it finds better solutions than the other two solvers in CVRP and CVRPL instances. These findings highlight the significant potential of heuristic algorithms in solving VRP variants effectively.
Table 4: Performance analysis of different solvers on three common VRP variants. The table presents the gaps compared to the results obtained by Gurobi. The best results among these solvers are highlighted in grey.
CVRP
Test sets N=25, C=200 N=50, C=200
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 223.6 18.24% 420.4 16.33%
Gurobi 189.1 0.00% 361.4 0.00%
Ours 186.9 -0.94% 358 -1.16%
CVRPTW
Test sets N=25, C=200 N=50, C=200
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 192.2 1.00% 392.0 8.47%
Gurobi 190.3 0.00% 361.4 0.00%
Ours 190.3 0.00% 361.4 0.00%
CVRPL
Test sets N=25, C=200, L=120 N=50, C=200, L=120
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 221.3 17.65% 393.4 9.04%
Gurobi 188.1 0.00% 360.8 0.00%
Ours 186.9 -0.64% 358 -0.78%
@@
5.3 Evaluation with Different LLMs
5.4 Ablation Study
An ablation study is conducted to evaluate the program development component of ARS. To analyze the impact of our database and constraint selection step on ARS, experiments are performed by removing each of these components separately.
As shown in Table 5, removing the constraint selection step leads to a decrease in the SR of ARS. This is because, without constraint selection, ARS acquires all six representative constraints, which are not necessarily all relevant to the current VRP. Consequently, this irrelevant information may mislead the LLM.
@@ 5.4 Ablation Study
Compared to constraint selection, removing the database has a more significant impact on the performance of ARS in generating correct programs. As previously discussed, ARS can be refined based on relevant constraints, thereby focusing on the distinctive aspects of a VRP variant. However, without reference to relevant constraints, the LLM must independently generate all the constraints, imposing a higher requirement on the LLM.
The results of this ablation study demonstrate that ARS achieves optimal program development performance only when both components are utilized.
6 Conclusion
In this paper, we have proposed ARS, a framework that leverages LLMs to automatically design constraint-aware heuristics to enhance a heuristic routing optimization framework for tackling VRP variants with complex and practical constraints. Moreover, we have developed and presented RoutBench, a comprehensive benchmark consisting of 1,000 VRP variants derived from 24 VRP constraints. Each variant includes a problem description, instance data, and validation code, enabling standard evaluation of routing solvers. The experimental evaluations of ARS have been particularly promising, demonstrating a superior performance over existing LLM-based methods and commonly used solvers. Specifically, ARS achieved a remarkable success rate of 91.7% in automatically solving common VRP instances and registered at least a 30% improvement across all tested benchmarks. In the future, we would like to extend the ARS framework to address other combinatorial optimization challenges, including scheduling and bin packing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Heuristics for VRPs
The Vehicle Routing Problem (VRP) is a classical combinatorial optimization problem that seeks optimal routes for vehicles to serve customers under constraints like capacity and time windows (Dantzig & Ramser, 1959). Variants such as the Capacitated VRP (CVRP) (Lysgaard et al., 2004), VRP with Time Windows (VRPTW) (Solomon, 1987), and Multi-Depot VRP (MDVRP) (Yuceer, 1997) have been extensively studied.
To address VRP’s computational challenges, heuristic methods are widely used. Local Search (LS) iteratively improves solutions using operators like 2-opt or 3-opt (Lin, 1965), while Iterated Local Search (ILS) introduces perturbations to escape local optima (Lourenço et al., 2003). Simulated Annealing (SA) explores worse solutions probabilistically to avoid stagnation (Kirkpatrick et al., 1983), and Tabu Search (TS) prevents revisiting solutions using a tabu list (Glover, 1989). Variable Neighborhood Search (VNS) systematically adjusts neighborhood structures for better exploration (Mladenović & Hansen, 1997).
@@
Adaptive Large Neighborhood Search (ALNS) selects neighborhood operators adaptively based on performance (Ropke & Pisinger, 2006). Recent ALNS variants integrate machine learning to predict effective operators (Mara et al., 2022) and reinforcement learning to optimize selection policies (Wang et al., 2024). Hybrid ALNS approaches, such as combining ALNS with branch-and-price, have further improved performance on constrained VRPs (Vidal, 2022).
In summary, these heuristic advancements have significantly promoted VRP research by enabling efficient solutions to increasingly complex and large-scale problem instances.
A.2 LLMs for VRPs
Recent advancements in large language models (LLMs) have opened new possibilities for solving vehicle routing problems (VRPs) by leveraging their ability to process and embed complex tasks through natural language understanding (Huang et al., 2024). These models have been applied to classical optimization problems such as the traveling salesman problem (TSP) and CVRP. In particular, LLMs, combined with evolutionary computation (EC), have demonstrated potential in automating the discovery of heuristics for VRPs (Liu et al., 2024a). For instance, guided local search and other algorithmic frameworks can be evolved with the assistance of LLMs, albeit requiring significant domain-specific expertise and computational resources for iterative model invocation (Wang et al., 2024).
@@ A.2 LLMs for VRPs
Another line of research emphasizes the transformation of textual problem descriptions into mathematical formulations and executable code that can be processed by external solvers (Tang et al., 2024). This approach benefits from LLMs’ ability to interpret user queries and generate structured outputs, enabling the automation of optimization tasks. Furthermore, multi-agent systems have been introduced to coordinate LLM-based agents for tasks such as problem formulation, programming, and evaluation (Xiao et al., 2023). While promising, these methods are inherently limited by the scope of the knowledge encoded within the pre-trained models, which restricts their applicability to novel or highly complex problems (Zhang et al., 2024).
To address these challenges, emerging studies focus on directly generating executable programs for VRPs by integrating LLMs with external knowledge sources. This approach bypasses the intermediate step of mathematical model formulation and instead leverages both the intrinsic reasoning capabilities of LLMs and external references to enhance problem-solving efficacy (Huang et al., 2024). Such integration aims to reduce reliance on domain-specific expertise while expanding the range of solvable problems, marking a shift towards more generalized and adaptive optimization frameworks (Romera-Paredes et al., 2024).
Appendix B Operator
Local search operators are critical components of heuristic methods, designed to explore the neighborhood of a solution and improve its quality. The commonly used operators are as follows:
2-opt Operator. The 2-opt operator is a classical approach originally developed for the Traveling Salesman Problem (TSP). It works by removing two non-adjacent edges in the solution and reconnecting them in a different way, thereby altering the order of nodes. If the new configuration reduces the total cost, it is accepted as an improved solution.
Swap Operator. The Swap operator exchanges the positions of two elements in the solution, generating a new configuration. This simple yet effective operator promotes diversity in the solution space and helps avoid local optima.
Shift Operator. The Shift operator moves an element from its current position to another position within the same solution or a different one. This adjustment redistributes the structure of the solution, enabling further exploration of the search space and improving solution quality.
Destroy Operator. The Destroy operator partially disrupts the current solution by selectively removing some elements, breaking its local optimality and enabling exploration of new regions in the search space. Two common implementations are Random Removal and String Removal:
  • Random Removal: This method removes elements uniformly at random, without relying on specific heuristics, making it a simple yet effective way to diversify the search process.
  • String Removal: This method targets sequences of consecutive or related elements (strings), such as partial routes or groups of customers (Christiaens & Vanden Berghe, 2020). It begins by randomly selecting a ”center” customer and removing a string of nearby customers from the route. The string size is randomly determined, constrained by the average route size and a predefined maximum. If constraints on the number of disrupted routes or previously disrupted routes are met, further removal is skipped.
Repair Operator. The Repair operator complements the Destroy operator by reinserting removed elements to reconstruct a complete solution, guided by optimization objectives. A commonly used implementation is the Greedy Repair:
  • Greedy Repair: This method reinserts removed elements one by one, selecting at each step the position that minimizes the objective function. By prioritizing optimal insertions, it balances solution quality with search efficiency.
These operators, through targeted adjustments, provide the foundation for heuristic and metaheuristic algorithms, enabling efficient exploration and exploitation of the solution space.
Appendix C Prompts of ARS

### v3

@@ 4.2 Analysis
The distribution of problems by the number of constraints is shown in Figure 3. The 48-problem subset consists of common problems, one is a simple VRP without any constraints, while the remaining 47 include one to five basic constraints. In RoutBench, the distribution reflects the proportions of the full set of 5624 problems. Specifically, problems with three constraints are the most common in the dataset, as three-constraint problems dominate the total number of problems with three or fewer constraints. For more complex problems, those with four and five constraints appear in similar proportions, ensuring a balanced representation of high-complexity scenarios.
@@
5 Experiments
To evaluate the performance of these LLM-based methods in handling different VRP variants, we assess their ability to generate correct programs within our solver. To analyze the impact of different solvers on the generated programs, we compare our solver with other solvers (e.g., CPLEX, OR-Tools, and Gurobi). Additionally, we included experiments with other open-source and closed-source LLMs (e.g., DeepSeek V3 and LLaMA 3.1 70B) to explore their impact on program generation. Finally, we also conduct an ablation study on our proposed ARS.
@@ 5 Experiments
Figure 4: Comparison of the success rate for different solvers with standard prompting.
Figure 5: Compares the lines of code generated by LLMs for different solvers on four VRP variants.
@@ 5.1 Comparison with LLM-Based Methods
The first experiment aims to evaluate the ability of ARS to produce successful programs. We compare our proposed ARS with seven LLM-based methods, including Standard Prompting, Chain of Thought (CoT) [60], Reflexion [61], Progressive-Hint Prompting (PHP) [62], Chain-of-Experts (CoE) [13], Self-debug [29], and Self-verification [12]. To focus on program generation, all methods are implemented using our backbone algorithm framework to handle various VRP variants.
This experiment is conducted on 48 common problems and RoutBench using GPT-4o. We use two metrics to evaluate the results. The Success Rate (SR) measures the proportion of the program where generated solutions pass the validation process. The Runtime Error Rate (RER) quantifies the percentage of programs that fail due to issues such as runtime errors, incorrect API usage, or syntax mistakes.
As shown in Table 2, ARS significantly outperforms other LLM-based methods in generating correct programs for both common problems and RoutBench. It is evident that these compared methods exhibit unsatisfactory performance across all problems, particularly on the RoutBench-H problems, where their SR is merely around 10%. None of the compared algorithms achieved an SR of 50% or higher. In contrast, ARS achieved an SR of 91.67% on the common problems. Moreover, on RoutBench, ARS generated correct solutions for 60% of the VRP variants, outperforming all seven other LLM-based methods by at least 30% in terms of SR. These results highlight ARS’s competitiveness in addressing complex VRP variants.
@@ 5.2 Comparison with Different Solvers
In the previous experiment, to focus on generating correct programs, all methods used the same solvers to handle VRP variants. To explore the impact of different solving tools on program generation, this experiment uses Standard Prompting with four solvers (CPLEX, OR-Tools, Gurobi, and our solver) for common problems.
As shown in Figures 5 and 5, under the standard prompting, our solver achieves the highest Success Rate (SR) on the common problems, and requires the fewest Lines of Code (LOC) compared to other solvers. This is because, in our framework, the LLM only needs to write constraint codes using straightforward Python syntax, while other solvers are limited by their specific implementations and require highly standardized modeling. For the same VRP variants, our solving framework makes program generation much simpler compared to these solvers. Examples of solver codes are provided in Appendix E.
To evaluate the performance of these solvers on different VRP variants, we conduct a comparative study with other solvers across four VRP variants (e.g., CVRP, CVRPTW, DCVRP, and DCVRP-L). Details of these problem sets are provided in Appendix D.6. These solvers usually require considerable time to solve VRPs, so we allocate 25, 50, and 100 minutes for problem sizes of 25, 50, and 100, respectively. As can be seen from Table 3, our solver achieved the optimal results in all instances. Although HGS and LKH-3 can solve common problems quickly and efficiently, our proposed method is designed to tackle a wide range of VRPs, even those it has never encountered before.
Table 3: Performance analysis of different solvers on four VRP variants. The table presents the gaps compared to the results obtained by ARS. The best results among the four solvers (e.g., CPLEX, OR-Tools, Gurobi, and ARS) are highlighted in grey.
CVRP CVRPTW
Solvers N=25 N=50 N=100 N=25 N=50 N=100
Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time
HGS 186.9 (0.00%) 0.1s 358.0 (0.00%) 0.2s 817.8 (0.00%) 0.3s 190.3 (0.00%) 0.1s 361.4 (0.00%) 0.1s 826.3 (0.00%) 0.4s
LKH-3 186.9 (0.00%) 0.1s 358.0 (0.00%) 0.1s 817.8 (0.00%) 1.3s 190.3 (0.00%) 0.1s 361.4 (0.00%) 0.1s 826.3 (0.00%) 4.2s
CPLEX 187.6 (0.37%) 45s 362.7 (1.31%) 47m 841.6 (2.91%) 1.1h 190.3 (0.00%) 1.3s 361.4 (0.00%) 5.2m 826.3 (0.00%) 1.3h
OR-Tools 186.9 (0.00%) 4.2s 358.8 (0.22%) 7.2s 849.2 (3.84%) 1.0m 190.3 (0.00%) 0.5s 362.5 (0.30%) 9.6s 828.1 (0.21%) 3.1m
Gurobi 186.9 (0.00%) 1.8m 358.0 (0.00%) 5.6m 828.0 (1.25%) 1.3h 190.3 (0.00%) 10s 361.4 (0.00%) 1.2m 826.3 (0.00%) 1.4h
Ours 186.9 (0.00%) 3.2s 358.0 (0.00%) 5.4s 817.8 (0.00%) 4.6m 190.3 (0.00%) 5.8s 361.4 (0.00%) 23s 826.3 (0.00%) 1.5m
DCVRP DCVRP-L
Solvers N=25 N=50 N=100 N=25 N=50 N=100
Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time Obj. $\downarrow$ (Gap) Time
CPLEX 213.3 (9.16%) 35s 392.1 (6.98%) 50m 878.4 (5.65%) 1.2h 215.3 (0.00%) 1.6m 390.7 (1.11%) 48m 871.0 (2.68%) 1.6h
OR-Tools 253.4 (29.68%) 0.7s 426.0 (16.23%) 39s 893.6 (7.48%) 1.4h 226.3 (5.11%) 1.4s 405.8 (5.02%) 1.5m 874.6 (3.10%) 11m
Gurobi 219.2 (12.18%) 5.2m 395.1 (7.80%) 40m 874.8 (5.22%) 1.1h 219.2 (1.81%) 2.4m 395.5 (2.36%) 44m 876.8 (3.36%) 1.0h
Ours 195.4 (0.00%) 6s 366.5 (0.00%) 1.8m 831.4 (0.00%) 18m 215.3 (0.00%) 2.4s 386.4 (0.00%) 26s 848.3 (0.00%) 19m
5.3 Evaluation with Different LLMs
@@
5.4 Ablation Study
An ablation study is conducted to evaluate the program generation component of ARS. To analyze the impact of our database and constraint selection step on ARS, experiments are performed by removing each of these components separately.
As shown in Table 4, removing the constraint selection step leads to a decrease in the SR of ARS. Without this step, ARS acquires all six representative constraints, which may not be relevant to the current VRP and can mislead the LLM. Compared to constraint selection, removing the database has a more significant impact on the performance of ARS in generating correct programs. As previously discussed, ARS can be refined based on relevant constraints, thereby focusing on the distinctive aspects of a VRP variant. However, without reference to relevant constraints, the LLM must independently generate all the constraints, imposing a higher requirement on the LLM.
The results of this ablation study demonstrate that ARS achieves optimal program generation performance only when both components are utilized.
6 Conclusion, Limitation, and Future Work
Conclusion. In this paper, we have proposed ARS, a framework that leverages LLMs to automatically design constraint-aware heuristics, enhancing a heuristic routing optimization framework for VRP variants with complex and practical constraints. Moreover, we have developed and presented RoutBench, a comprehensive benchmark consisting of 1,000 VRP variants derived from 24 VRP constraints. Each variant includes a problem description, instance data, and validation code, enabling standard evaluation of routing solvers. The experimental evaluations of ARS have been particularly promising, demonstrating a superior performance over existing LLM-based methods and commonly used solvers.
Limitation and Future Work. Designing specialized algorithms for various problems often achieves better solving capability and efficiency. In this paper, we mainly focus on generating correct solving programs for a wide range of VRPs using a unified heuristic solving framework. In the future, the solver framework can be enhanced to incorporate LLMs for automated algorithm design tailored to different problems, further improving its solving capability and efficiency.
@@ Appendix A Related works
The Vehicle Routing Problem (VRP) is a classical combinatorial optimization problem that seeks optimal routes for vehicles to serve customers under constraints like capacity and time windows  [63]. Over the years, many VRP variants have been developed and extensively studied, including the Capacitated VRP (CVRP)  [64], the VRP with Time Windows (VRPTW)  [15], and the Multi-Depot VRP (MDVRP)  [17]. However, solving these VRP variants often requires experts to deeply understand the specific problem characteristics, including the constraints, customer demands, and operational rules. This process involves carefully analyzing the problem, designing appropriate models, and implementing customized algorithms. Such methods usually require complex coding and are limited in their ability to address only a small number of VRP variants, making them less flexible and scalable for the diverse and evolving challenges in real-world applications.
@@ A.1 Heuristics for VRPs
Traditional methods for solving VRP often rely on heuristics. Some simple heuristics, such as the Greedy algorithm and hill-climbing, are commonly used to solve VRP. The Greedy algorithm builds a solution step-by-step by making the most immediate, optimal choice at each step, though it often leads to suboptimal global solutions [65]. Hill-climbing, on the other hand, iteratively improves a solution by moving to a better neighboring solution, but it is prone to getting stuck in local optima [66]. To address these limitations, more advanced metaheuristics have been developed. Simulated Annealing (SA) probabilistically accepts worse solutions during the search process to escape local optima, mimicking the physical annealing process [67]. Tabu Search (TS) enhances local search by using a tabu list to prevent revisiting previously explored solutions, enabling it to explore broader solution spaces [68].
In addition to these, state-of-the-art approaches like the Hybrid Genetic Search (HGS) and Lin-Kernighan-Helsgaun (LKH-3) algorithms have achieved remarkable success in solving VRPs. HGS combines genetic algorithms with heuristics tailored for specific types of VRPs, efficiently balancing exploration and exploitation [4]. It is particularly powerful for large-scale and complex VRPs. LKH-3, an extension of the classic Lin-Kernighan heuristic, is highly effective for solving Traveling Salesman Problems (TSP) and TSP-based VRPs, leveraging advanced search strategies and efficient implementations to achieve near-optimal solutions [69].
@@
Adaptive Large Neighborhood Search (ALNS) represents a more dynamic and flexible approach [70]. It adaptively selects different neighborhood operators based on their performance during the search process, making it highly effective for solving complex VRP variants [71]. Recent advancements in ALNS integrate machine learning techniques to predict the most effective operators and reinforcement learning to optimize selection policies [72]. Hybrid ALNS approaches, such as combining ALNS with branch-and-price methods, further enhance their ability to solve constrained and large-scale VRPs [30].
Despite their success, these methods often require experts to deeply understand the specific VRP variant, carefully model the problem, and implement customized algorithms. This reliance on expert knowledge limits the scalability of these approaches to handle a broader range of VRP variants without significant manual effort.
A.2 NCO for VRPs
Neural combinatorial optimization (NCO) represents a paradigm shift in solving VRPs by leveraging deep learning models to directly learn problem-solving strategies from data. Several notable NCO approaches have been proposed to address the challenge of solving multiple VRP variants within a unified framework. For instance, MTPOMO tackles cross-problem generalization by representing VRPs as combinations of shared attributes, allowing a single model to solve unseen variants in a zero-shot manner [73]. MVMoE employs a multi-task learning framework with a mixture-of-experts architecture, using a hierarchical gating mechanism to balance model capacity and computational efficiency, achieving strong results across ten unseen VRP variants [74]. CaDA further advances the field by incorporating a constraint-aware dual-attention mechanism, which effectively captures both global and local problem-specific information, enabling state-of-the-art performance on sixteen VRP variants [75].
However, despite these advancements, current NCO methods still face significant limitations. They typically require manual modifications to adapt algorithms for new VRP variants, limiting their scalability and practicality for real-world applications with highly diverse constraints.
A.3 LLMs for VRPs
Recent advancements in large language models (LLMs) have introduced new possibilities for solving vehicle routing problems (VRPs) by leveraging their capacity to encode and process complex optimization tasks [12]. LLM-based automatic heuristic design (AHD) has emerged as a promising approach, enabling the generation of high-quality heuristics for problems like the traveling salesman problem (TSP) and capacitated VRP (CVRP) without extensive domain expertise. Methods such as Evolutionary Optimization Heuristics (EoH) integrate LLMs with evolutionary computation (EC) to iteratively refine a population of heuristics, facilitating automated discovery of effective solutions [76]. However, population-based approaches often converge prematurely to local optima. To overcome this, Monte Carlo Tree Search-based AHD (MCTS-AHD) organizes LLM-generated heuristics into a tree structure, enabling deeper exploration of the search space and better utilization of underperforming heuristics [77].
Other studies have explored the application of LLMs to different VRP variants, showcasing innovative approaches and promising results. For example, LLM-driven Evolutionary Algorithms (LMEA) utilize LLMs as evolutionary optimizers, achieving competitive results on TSPs with minimal domain knowledge [78]. Mechanisms like self-adaptation help balance exploration and exploitation, effectively avoiding local optima. Similarly, an approach proposed by Huang et al. [12] enables LLMs to directly generate executable programs for VRPs from natural language task descriptions. This method is further enhanced by a self-reflection framework, which allows LLMs to debug and verify their solutions, significantly improving feasibility, optimality, and efficiency. These early explorations highlight the potential of LLMs in addressing VRPs and advancing the field.
@@ A.3 LLMs for VRPs
Another line of research explores transforming textual problem descriptions into mathematical formulations and executable code that can be processed by external solvers [79]. This approach benefits from LLMs’ ability to interpret user queries and generate structured outputs, enabling the automation of optimization tasks. In parallel, multi-agent systems have been introduced to coordinate LLM-based agents for tasks such as problem formulation, programming, and evaluation [13]. Separately, the DRoC framework introduces a novel method for solving complex VRPs by decomposing constraints, retrieving external knowledge through a retrieval-augmented generation (RAG) approach, and integrating it with the model’s internal knowledge. By dynamically optimizing program generation, this framework has demonstrated significant improvements in both accuracy and runtime efficiency across 48 VRP variants [20]. However, despite these innovations, these methods remain inherently constrained by the scope of knowledge encoded within pre-trained models, particularly in generating solver-specific code. This limitation poses significant challenges for LLMs in addressing novel or highly complex problems [14].
Appendix B Detailed Methodology
B.1 Prompts of ARS